OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5 They didn't teach this at all, the lesson was all about verifying and I'm so stuck on what to replace sin^2 and sin with to actually get anywhere

OpenStudy (gorv):

7sin^2x-14sinx+7=0 sin^2x-2sinx+1=0 (sinx-1)^2=0 sinx-1=0 sinx=1....x=pi/2 x=multiples of pi/2

OpenStudy (gorv):

@heretolearnstuff

OpenStudy (anonymous):

Thanks! I actually figured it out (I'm used to not getting responses) But thank you for taking the time to respond @gorv I appreciate it :)

OpenStudy (gorv):

always welcome buddy @heretolearnstuff